# Project page: https://zhang-tao-whu.github.io/projects/DVIS_DAQ DVIS-DAQ: Improving Video Segmentation via Dynamic Anchor Queries

### v1

@@ 3.2 Dynamic Anchor Queries
Dynamic Anchor Queries for Emergence. When processing the Tth frame of a video, we consider all objects appearing in the current frame as candidates for newly emerging objects. The dynamic anchor queries for emergence DAQEmg are naturally generated based on the features of these candidates. As shown in Fig. 3 (a), DAQEmg consists of two parts: query feature DAQ and query positional embedding DAQ. The query feature DAQ is a learnable embedding shared by all dynamic anchor queries. The query positional embedding DAQ is obtained using the appearance features of the candidates through simple Mask-Pooling:
@@
Dynamic Anchor Queries for Disappearance. When processing the Tth frame, we consider all currently tracked objects as candidates for disappearing objects. Similar to the dynamic anchor queries for emergence, the DAQDis for disappearing objects are naturally generated based on the features of these candidates. As shown in Fig. 3 (b), the DAQDis also consists of query feature DAQ and query positional embedding DAQ. We use the candidates’ momentum-weighted appearance features as the query positional embedding DAQ. The momentum-weighted function follows CTVIS [72]:
@@
Continuously Tracked Queries. We also assign query positional embedding CTQpos for continuously tracked queries (CTQ) to maintain the format consistency with dynamic anchor queries. We use the momentum-weighted appearance feature of the tracked objects as CTQpos, the same process as Eq. 3, 4, and 5.
Incorporating Dynamic Anchor Queries into a Tracker. As depicted in Fig. 4, we can seamlessly integrate DAQ into mainstream propagation-based trackers [75, 24], requiring no alterations to the tracker’s network architecture. We instantiate two trackers: Tracker 1, which takes CTQ and DAQEmg as inputs and is responsible for tracking continuously appearing and newly emerging objects, and Tracker 2, which receives DAQDis along with several learnable background embeddings as inputs and identifies disappearing objects in the current frame. It’s important to note that the SoftMax function in Tracker 2 operates along the Q-dimension to prevent multiple dynamic anchor queries from segmenting the same object. Through this pipeline, we can effectively and reasonably manage and model the emergence and disappearance of objects. Despite introducing an additional tracker, it’s worth highlighting that the computational overhead is minimal, constituting less than 2% of the overall computational cost [75], as processing occurs solely at the query level rather than on dense image features.
@@ 3.3 Emergence and Disappearance Simulation
New Emergence Simulation. As shown in Fig. 5 (a), taking the green object as an example, we can simulate it as an emerging object by removing its corresponding query from CTQ. At this point, this green object is not present in CTQ but exists in the current frame, consistent with the state of a real newly emerged object colored in gray. Then, the Tracker needs to transition the corresponding anchor query from DAQEmg to a foreground query and predict this green object’s segmentation mask and class.
@@ 4.2 Main Experiments
Performance on Youtube-VIS 2022 dataset. Youtube-VIS 2022 has been expanded with 71 long videos based on the 2021 version. Therefore, we only report the APL for the long videos in Tab. 2. When using ResNet-50 as the backbone, DVIS-DAQ outperforms the baseline DVIS with 3.0 APL. Our method outperforms DVIS++ with 4.5 APL (42.0 vs. 37.5) when using VIT-L as the backbone.
@@ 4.3 Ablation Studies and Analysis
Qualitative results comparison. In Fig. 7, we present the visual comparison with our baseline. In this challenging scene, there are new-emerging, disappearing, and reappearing video objects. Our method accurately addresses all these scenarios, whereas DVIS suffers from issues including missed new-emerging objects and ID switches. Although DVIS detected object 1 in frame 2, it incorrectly switched the ID of another object to that of object 1. In the case of the reappearance of objects 1 and 2 in Frame 6, our method successfully maintained continuous tracking. In Frame 7, DVIS incorrectly switched the IDs of two nearby disappearing objects to objects 3 and 4, whereas our method correctly handled these disappearing cases.
@@ 0.A.4 More Qualitative Results
| GenVIS‡ [24] | Swin-L | 63.8 | 85.7 | 68.5 | 56.3 | 68.4 | 60.1 | 80.9 | 66.5 | 49.1 | 64.7 |
| MDQE‡ [32] | Swin-L | 59.9 | - | - | - | - | 56.2 | 80.0 | 61.1 | 44.9 | 59.1 |
| NOVIS‡ [41] | Swin-L | 65.7 | 87.8 | 72.2 | 56.3 | 70.3 | 59.8 | 82.0 | 66.5 | 47.9 | 64.6 |
| Tube-Link‡ [37] | Swin-L | 64.6 | 86.6 | 71.3 | 55.9 | 69.1 | 58.4 | 79.4 | 64.3 | 47.5 | 63.6 |
| RefineVIS‡ [1] | Swin-L | 64.4 | 88.3 | 72.2 | 55.8 | 68.4 | 61.2 | 83.7 | 69.2 | 47.9 | 64.8 |

### v2

@@ 3.2 Dynamic Anchor Queries
Dynamic Anchor Queries for Emergence. When processing the T frame of a video, we consider all objects appearing in the current frame as candidates for newly emerging objects. The dynamic anchor queries for emergence DAQ are naturally generated based on the features of these candidates. As shown in Fig. 3 (a), DAQ consists of two parts: query feature DAQ and query positional embedding DAQ. The query feature DAQ is a learnable embedding shared by all dynamic anchor queries. The query positional embedding DAQ is obtained using the appearance features of the candidates through simple Mask-Pooling:
@@
Dynamic Anchor Queries for Disappearance. When processing the T frame, we consider all currently tracked objects as candidates for disappearing objects. Similar to the dynamic anchor queries for emergence, the DAQ for disappearing objects are naturally generated based on the features of these candidates. As shown in Fig. 3 (b), the DAQ also consists of query feature DAQ and query positional embedding DAQ. We use the candidates’ momentum-weighted appearance features as the query positional embedding DAQ. The momentum-weighted function follows CTVIS [72]:
@@
Continuously Tracked Queries. We also assign query positional embedding CTQ for continuously tracked queries (CTQ) to maintain the format consistency with dynamic anchor queries. We use the momentum-weighted appearance feature of the tracked objects as CTQ, the same process as Eq. 3, 4, and 5.
Incorporating Dynamic Anchor Queries into a Tracker. As depicted in Fig. 4, we can seamlessly integrate DAQ into mainstream propagation-based trackers [75, 24], requiring no alterations to the tracker’s network architecture. We instantiate two trackers: Tracker 1, which takes CTQ and DAQ as inputs and is responsible for tracking continuously appearing and newly emerging objects, and Tracker 2, which receives DAQ along with several learnable background embeddings as inputs and identifies disappearing objects in the current frame. It’s important to note that the SoftMax function in Tracker 2 operates along the Q-dimension to prevent multiple dynamic anchor queries from segmenting the same object. Through this pipeline, we can effectively and reasonably manage and model the emergence and disappearance of objects. Despite introducing an additional tracker, it’s worth highlighting that the computational overhead is minimal, constituting less than 2% of the overall computational cost [75], as processing occurs solely at the query level rather than on dense image features.
@@ 3.3 Emergence and Disappearance Simulation
New Emergence Simulation. As shown in Fig. 5 (a), taking the green object as an example, we can simulate it as an emerging object by removing its corresponding query from CTQ. At this point, this green object is not present in CTQ but exists in the current frame, consistent with the state of a real newly emerged object colored in gray. Then, the Tracker needs to transition the corresponding anchor query from DAQ to a foreground query and predict this green object’s segmentation mask and class.
@@ 4.2 Main Experiments
Performance on Youtube-VIS 2022 dataset. Youtube-VIS 2022 has been expanded with 71 long videos based on the 2021 version. Therefore, we only report the AP for the long videos in Tab. 2. When using ResNet-50 as the backbone, DVIS-DAQ outperforms the baseline DVIS with 3.0 AP. Our method outperforms DVIS++ with 4.5 AP (42.0 vs. 37.5) when using VIT-L as the backbone.
@@ 4.3 Ablation Studies and Analysis
Qualitative results comparison. In Fig. 7, we present the visual comparison with our baseline. In this challenging scene, there are new-emerging, disappearing, and reappearing video objects. Our method accurately addresses all these scenarios, whereas DVIS suffers from issues including missed new-emerging objects and ID switches. Although DVIS detected object 1 in frame 2, it incorrectly switched the ID of another object to that of object 1. In the case of the reappearance of objects 1 and 2 in Frame 6, our method successfully maintained continuous tracking. In Frame 7, DVIS incorrectly switched the IDs of two nearby disappearing objects to objects 5 and 6, whereas our method correctly handled these disappearing cases.
@@ 0.A.4 More Qualitative Results
| GenVIS [24] | Swin-L | 63.8 | 85.7 | 68.5 | 56.3 | 68.4 | 60.1 | 80.9 | 66.5 | 49.1 | 64.7 |
| MDQE [32] | Swin-L | 59.9 | - | - | - | - | 56.2 | 80.0 | 61.1 | 44.9 | 59.1 |
| NOVIS [41] | Swin-L | 65.7 | 87.8 | 72.2 | 56.3 | 70.3 | 59.8 | 82.0 | 66.5 | 47.9 | 64.6 |
| Tube-Link [37] | Swin-L | 64.6 | 86.6 | 71.3 | 55.9 | 69.1 | 58.4 | 79.4 | 64.3 | 47.5 | 63.6 |
| RefineVIS [1] | Swin-L | 64.4 | 88.3 | 72.2 | 55.8 | 68.4 | 61.2 | 83.7 | 69.2 | 47.9 | 64.8 |